a car travels 150 km on 12L of gasoline.how many liters of gasoline are needed to travel 1500km?
Answer by checkley71(8403)

150/12=1500/X CROSS MULTIPLY
150X=12*1500
150X=18.000
X=18,000/150
X=120 LITERS OF GAS WILL BE NEEDED TO TRAVEL 1500 KM.